Fermi National Laboratory


Accelerator Update - June 13-June 16

Accelerator Overview

Friday June 13
The day shift began with repairs to A:LQ, so no stacking was taking place. The Tevatron and SY120 were doing studies and MiniBooNe was taking beam.

At 9:25 AM the repairs were complete on A:LQ and we were stacking by 10:12 AM.

At ~12:12 PM SY120 studies were complete for the day. Beam was delivered to the absorber. There are a few instrumentation repairs to complete next week.

At about 12:27 PM, a power glitch caused some of the accelerator systems to trip off.

Feeder 44, which feeds Wilson Hall, MiniBooNe, NuMI, the Leaderman Center, and some other areas, tripped off due to a cable fault. Within a short period of time, power was rerouted and restored to these areas. The cable fault will be investigated and repaired at a later date.

Operations started to recover systems.

By 1:47 PM, Linac was back up and running and beam was back to Booster by 2:16 PM.

After the TeV was recovered, the D0 and CDF experiments made controlled accesses.

Pbar resumed stacking at 3:30 PM.

Booster tuning and problems with the TEL brought us into the evening shift. D0 completed their access earlier and CDF was out by 6:15 PM.

By 10:30 PM, the TeV was ramping and doing studies, the TEL was repaired and things were looking good.

Saturday June 14
The midnight shift began with the TeV working on orbit studies, and with MiniBooNe taking beam. The plan was to start a shot setup after studies.

At 3:50 AM, shot setup began.

A store with the initial luminosity of 30.0E30 was in the TeV by 6:00 AM.

The remainder of the shift was spent stacking, running beam to MiniBooNe and with a store in the TeV.

The store was terminated around 9:00 PM, and both CDF and D0 were allowed to make accesses. By 11:00 PM, they were out and the TeV was ramping again.

Shot setup was under way by 11:30 PM.

Sunday June 15
The midnight shift began with Operations working on the shot and MiniBooNe taking beam.

Operations had the store in the TeV with an initial luminosity of 34.4E30 by 2:00 AM. Pbar returned to stacking

The store stayed in and stacking continued throughout the day. The store was terminated at 6:00 PM and shot setup began within the hour.

There were problems during the shot and some protons had to be reloaded. There were also problems with some of the Pbar shots.

Despite the problems, the store was in the TeV by 9:00 PM. It had an initial luminosity of 28.0E30.

Monday June 16
The midnight shift began with the TeV in a store, Pbar stacking, and with MiniBooNe taking beam.

Stacking resumed although there were some Linac trips and Booster vacuum problems.

Today's Plan
The Run Coordinator says there will be no studies this week, but an access into the Main Injector tomorrow for Recycler work. The store will be terminated later this morning and another store will be put in the TeV. The TeV will have a store tomorrow and will not be accessed during the MI access time.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/19/2003   email Fermilab